(Mathematical statistics)

* If Z1 , Z_{2} are independent standard normal random variables,

find the density of \frac{Z_{1}}{Z_{2}}.
Answer #1

i Given. 21 22 be independent standard normal (21) -L2, co f (21) = 21 -BL2ILOS 2n [ A(z) - e z ? - <Z2 LA 27 joind density of(YY) = -(4,²4² +42 ²/2 27 -OLY, LOS fluu42) o e-42 (41²+1) 12 -OLY2LOO 2T 22 We have to find dencity of 21 - 4 - f(1) = f( 2RIA) = + en ot o acould ; -olyi Lao 77 (4+) Cauchy distribution, which is odencity of standard i e fly = f(21) occoil).